40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 25 23:15:11 2013, Masayuki, Summary, Green Locking, FPMI noise caused by ARM locking fpmibd.pdf
    Reply  Wed Sep 25 23:59:29 2013, Masayuki, Summary, Green Locking, FPMI noise caused by ARM locking XARM_OLTF.pdfYARM_OLTF.pdfXARM_ITMXresponse.pngYARM_ITMYresponse.pngfree_running.pdf
       Reply  Thu Sep 26 01:49:28 2013, Masayuki, Summary, Green Locking, FPMI noise caused by ARM locking MICH_OLTF.pdfBS-RS55Q.pngETMX-RS55.pngETMY-RS55.pngplot.pdf
       Reply  Thu Sep 26 23:02:40 2013, rana, Update, LSC, FPMI noise caused by ARM locking Xarm_sweep_130926.pdflsc.pdferr.png
          Reply  Fri Sep 27 00:48:53 2013, Masayuki, Update, LSC, FPMI noise caused by ARM locking phase_badget.pngFPMI_noises.png
Message ID: 9162     Entry time: Wed Sep 25 23:59:29 2013     In reply to: 9161     Reply to this: 9163   9167
Author: Masayuki 
Type: Summary 
Category: Green Locking 
Subject: FPMI noise caused by ARM locking 

 

Measurement with ARMs

i) By locking the MICH with AS55Q signal I measured the actuator response of ITMX ITMY BS for calibration of each actuator. This measurement was done at the same time with elog#9158. The actuator response was
 
BS : 2.2347e-8 / f^2 [m/count]

ITMX: 5.0843e-9 /f^2 [m/count]

ITMY: 4.9677e-9 / f^2 [m/count]

 
 
ii)By locking the Arms for IR with POX,POY. I measured the OLTF and the response from ITM actuation to POX and POY signal. Attachment 1,2 are the plots of fitted OLTF , the measured OLTF, and residual function (model - measure)/model and the attachment 3,4 are the response of each arm. I fitted the three parameters. Those are the gain, time-delay and cavitypole. Each fitted parameter is
 
XARM ;
timedelay:-282.09 usec, cavity pole : 2872.0 Hz
YARM ;
timedelay:-283.84 usec, cavity pole : 2939.9 Hz
 
The cavity pole seems higher than privious measurement (In 2009). Actually the residual function are increase at the higher frequency region than 1kHz, so I guess the fitting is not so good.One possibility is that in the region where cavity pole effect increase we has not much data.
With fitted OLTF and actuator response I calibrated the H_xarm and H_yarm.
 
Hxarm : 2.9796 e11 [count / m]
Hyarm : 6.1394 e11 [count / m]
 
iii) After that I measured the response from ETM actuation to POX and POY signal to calibrate the ETM actuator. The response of each actuator is
 
ETMX:1.2040e-8 / f^2 [m/count]

ETMY:1.4262e-8 / f^2 [m/count]
 
iv) I calibrated the error signal with OLTF and Hxarm,Hyarm. The result is in Attachment 5

 In high frequency region there is the difference between xarm and yarm. These difference are already there in error signal. I'm not sure where these noise comes from. We will make measurement with Green PDH from tomorrow, so  we can also check with those measurement.

In other region the two noises are very close and also very similar to the plot of the seismic motion in the control room (attached on the front of TV screen).

Attachment 1: XARM_OLTF.pdf  42 kB  | Hide | Hide all
XARM_OLTF.pdf
Attachment 2: YARM_OLTF.pdf  43 kB  | Hide | Hide all
YARM_OLTF.pdf
Attachment 3: XARM_ITMXresponse.png  72 kB  | Hide | Hide all
XARM_ITMXresponse.png
Attachment 4: YARM_ITMYresponse.png  72 kB  | Hide | Hide all
YARM_ITMYresponse.png
Attachment 5: free_running.pdf  76 kB  Uploaded Thu Sep 26 01:30:09 2013  | Hide | Hide all
free_running.pdf
ELOG V3.1.3-